You have 2 positive numbers. One number is one-fifth of the other number. The difference between the two numbers is 284, find the numbers.

Found 2 solutions by math_tutor2020, ikleyn:
Answer by math_tutor2020(3817)


x = smaller number
5x = larger number

The ratio of the two values is
smaller/larger = (x)/(5x) = 1/5
Which shows that the smaller is 1/5 of the larger
Flipped around: the larger is 5 times that of the smaller

The difference in the numbers is 284
larger - smaller = 284
5x - x = 284
4x = 284
x = 284/4
x = 71
So,
5x = 5*71 = 355

Answer:
The numbers are 71 and 355


Answer by ikleyn(52847)

.
You have 2 positive numbers. One number is one-fifth of the other number.
The difference between the two numbers is 284, find the numbers.
~~~~~~~~~~~~~~~~~

        Algebra solution



The numbers are  x  and  5x.

Their difference is  284:

        5x - x = 284

          4x     = 284

            x     = 284/4 = 71.

ANSWER.     The numbers are  71  and  5*71 = 355.

Solved.

        Mental solution



The difference of  284  is  4/5  of the greater number.

Hence,  1/5  of the greater number is  1/4  of  284,  or  71.

It means that the greater number is  355  and the smaller number is  355:5 = 71.

You get the same answer.


                Solved twice . . .


You may choose that solution which you like more . . .